The relatively high resistivity of dry skin, about 1 × 106 O·m, can safely limit the flow of current into deeper tissues of the body. Suppose an electrical worker places his palm on an instrument whose metal case is accidentally connected to a high voltage. The skin of the palm is about 1.5 mm thick. Estimate the area of skin on the worker’s palm that would contact a flat panel, then calculate the approximate resistance of the skin of the palm.
Answer #1

Area of the flat palm is \(A=\) Length times breadth

$$ \begin{array}{l} =(10 \mathrm{~cm})(15 \mathrm{~cm}) \\ =150 \mathrm{~cm}^{2} \\ =150 \mathrm{~cm}^{2}\left(\frac{10^{-4} \mathrm{~m}^{2}}{1 \mathrm{~cm}^{2}}\right) \\ =1.5 \times 10^{-2} \mathrm{~m}^{2} \end{array} $$

Thickness of skin palm is \(l=1.5 \mathrm{~mm}\)

$$ \begin{aligned} \text { Resistance } R &=\frac{\rho l}{A}=\frac{\left(1.0 \times 10^{6} \Omega \cdot \mathrm{m}\right)\left(1.5 \times 10^{-3} \mathrm{~m}\right)}{1.5 \times 10^{-2} \mathrm{~m}^{2}} \\ &=1.0 \times 10^{5} \Omega \end{aligned} $$